OpenStudy (anonymous):

3(2k-2)=-2(9k-11) please show work.

OpenStudy (anonymous):

6k-6=-18k+22 6k+18k=22+6 24k=28 k=28/24 k=7/6

OpenStudy (anonymous):

the answer is supposed to be 2 though?